# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v1

@@ 5 Experiments
In the following sections, we investigate whether combining LRRL with Adam or RMSProp —two widely used optimizers— can improve cumulative returns in deep RL algorithms. To assess this, we compare LRRL against learning methods with and without schedulers using the baseline implementation of DQN provided in Dopamine [9]. We test LRRL under different configurations and Atari games, reporting the average and standard deviation of returns over 5 runs. Details on the evaluation metrics and hyperparameters used in these experiments are summarized in Appendix B.
@@ 5.1 Comparing LRRL with Standard Learning
In our first experiment, we consider a set of $5$ learning rates, and compare the performance of 5 configurations of LRRL, each of them using a subset of those learning rates, against the DQN algorithm reaching best performance, in terms of maximum average return, among the $5$ possible learning rates choices (see Figure 5 in Appendix A.1). More precisely, the set of learning rates is
@@
The whole set $\mathcal{K}(5)$ is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) and 3 taking the lowest/middle/highest ($\mathcal{K}_{\text{sparse}}(3)$) learning rate values. All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations.
Results are gathered in Table 1 and illustrated in Figure 1 along with four Atari games. They show that LRRL outperforms standard learning in two out of four tasks while remaining competitive in the others. Notably, LRRL not only matches or exceeds performance but also reduces the need for extensive parameter tuning by incorporating multiple learning rates in a single run. However, as the results indicate, the $3$-arm bandit variants exhibit different behaviors during the learning process, suggesting that the choice of the number of arms and their values still requires task-specific tuning to achieve good performance.
Figure 1: A comparison between 5 configurations of LRRL with various learning rate subsets and the DQN algorithm reaching the best performance among possible learning rates.
Game DQN LRRL $\mathcal{K}_{\text{lowest}}(3)$ LRRL $\mathcal{K}_{\text{middle}}(3)$ LRRL $\mathcal{K}_{\text{highest}}(3)$ LRRL $\mathcal{K}_{\text{sparse}}(3)$ LRRL $\mathcal{K}(5)$
Asteroids 1 085 $\pm$ 63 1 065 $\pm$ 70 1 079 $\pm$ 48 1 061 $\pm$ 59 1007 $\pm$ 127 1 085 $\pm$ 88
Breakout 217 $\pm$ 14 220 $\pm$ 24 201 $\pm$ 6 177 $\pm$ 8 270 $\pm$ 13 253 $\pm$ 20
Pong 19 $\pm$ 0.4 18 $\pm$ 0.3 19 $\pm$ 0.9 19 $\pm$ 0.5 19 $\pm$ 0.5 19 $\pm$ 0.8
Seaquest 5 881 $\pm$ 1 533 5 905 $\pm$ 2 557 6 135 $\pm$ 2 229 6 231 $\pm$ 1 802 8 920 $\pm$ 2 759 6 799 $\pm$ 2 060
Table 1: Max average return (best in bold if significantly better than others) and its standard deviation for 4 Atari games.
To illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves, Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$. In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses. The exception is in the Pong environment, where the model converges after only a few iterations, resulting in a more uniform probability distribution across the set of arms.
Figure 2: Systematic sampling of normalized learning rates and returns over the training steps using LRRL $\mathcal{K}(5)$ with Adam optimizer, through a single run. For each episode, we show the selected learning rate using different colors.
@@ 5.2 Combining and Comparing Schedulers with LRRL
Next, we consider using LRRL combined with learning rate schedulers. Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds). We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
@@
Figure 3 and Table 2 show that LRRL combined with schedulers can substantially increase final performance compared to using exponential decay schedulers for some environments while remaining competitive for others. The dashed black line represents the max average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by [4], who linearly decay the learning rate to 00.
Figure 3: A comparison between LRRL with arms as schedulers and schedulers individually.
Game DQN $d=1\times 10^{-7}$ $d=2\times 10^{-7}$ $d=3\times 10^{-7}$ LRRL ($\mathcal{K}_{s}$)
Asteroids 1 028 $\pm$ 53 1 025 $\pm$ 52 1 013 $\pm$ 56 1 063 $\pm$ 82 1 015 $\pm$ 33
Breakout 144 $\pm$ 12 149 $\pm$ 11 151 $\pm$ 7 144 $\pm$ 12 233 $\pm$ 19
Seaquest 5 881 $\pm$ 1 533 5 284 $\pm$ 1 134 5 793$\pm$ 1 225 6 612 $\pm$ 835 13 864 $\pm$ 3 581
Video Pinball 410 186 $\pm$ 193 328 388 768 $\pm$ 195 150 450 015 $\pm$ 178 876 362 645 $\pm$ 172 169 388 308 $\pm$ 103 862
Table 2: Max average return (best in bold) and its standard deviation for 4 Atari games.
5.3 RMSProp Optimizer and more Environments
@@
Another widely used optimizer for training deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism. Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum. Although the standard RMSProp does not feature momentum, we found that adding momentum to RMSProp can increase both the performance of DQN and LRRL, aligning with findings in the literature [28, 4].
@@
In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN. As shown in Figure 4 and Table 3, LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN. Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance. Future work should investigate whether this slow convergence is linked to factors such as the environment’s stochasticity or the optimizer’s features such as the absence of bias correction in the first and second moment estimates.
@@
Game DQN (RMSProp-M) LRRL (RMSProp-M) DQN (Adam) LRRL (Adam)
Asterix 11 464 $\pm$ 2 848 6 499 $\pm$ 927 12 561 $\pm$ 1 245 15 017 $\pm$ 3 892
Ms. Pacman 3 301 $\pm$ 310 2 696 $\pm$ 248 3 232 $\pm$ 114 3 310 $\pm$ 231
Space Invaders 1 490 $\pm$ 132 2 712 $\pm$ 73 2 874 $\pm$ 319 3 641 $\pm$ 1 030
Table 3: Max average return (best in bold) and its standard deviation for 3 Atari games.
@@ 6 Conclusion
In this work, we introduced dynamic Learning Rate for Deep Reinforcement Learning (LRRL), a meta-learning approach for selecting the optimizer’s learning rate on the fly. We demonstrated empirically that combining LRRL with the Adam optimizer could significantly enhance the performance of the value-based algorithm DQN, outperforming baselines and learning rate schedulers in some tasks while remaining competitive in others. Furthermore, by employing a multi-armed bandit algorithm, LRRL reduces the need for extensive hyperparameter tuning, as it explores a set of learning rates in a single run with minimal extra computational overhead.
While this work focused on dynamically selecting the best-performing learning rate, future investigations could extend LRRL ideas to other critical hyperparameters, such as mini-batch size, which also plays a key role in the model’s convergence. Moreover, although LRRL selects learning rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [30, 18, 1].
@@ Acknowledgment
Appendix A Supplementary Experiments
A.1 Baseline Evaluation with Varying Learning Rates
@@
To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer. The results presented in Figure 5 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to the worst jumpstart performance and slow convergence.
@@
Figure 5: DQN performance using Adam optimizer with varying learning rates across 4 Atari games.
A.2 A Comparison between Multi-Armed Bandits Algorithms
@@
Adversarial MAB algorithms are designed for environments where the reward distribution changes over time. In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown probability distributions. To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5]. MOSS trade-off exploration-exploitation by pulling the arm $k$ with highest upper confidence bound given by:
@@
where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to timestep $t$.
In Figure 6, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balance exploration and exploitation in MOSS. Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$. The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
Figure 6: A comparison between adversarial and stochastic MAB algorithms.
Appendix B Experiments Details
B.1 Evaluation Terminology
In this section, we describe the evaluation metrics that can be used to evaluate agent’s performance as it interacts with an environment. Based on Dopamine, we use the evaluation step-size “iterations”, which is defined as a predetermined number of episodes. Figure 7 illustrates the evaluation metrics used in this work, as defined in [36]:

### v2

@@ 4 Experiments
In the following sections, we examine whether combining LRRL with Adam or RMSProp —two widely used optimization algorithms— can enhance cumulative returns in deep RL algorithms. We compare LRRL against learning methods both with and without schedulers, using baseline algorithms provided in the Dopamine [9] framework. The experiments involve testing LRRL across various configurations and a selection of Atari games, reporting the mean and one-half standard deviation of returns over 5 independent runs. As testing the full Atari benchmark is extremely costly, we choose a representative subset of games guided by their baseline learning curves, which offer insights into reward sparsity, stochasticity, and exploration demands. Further details on the evaluation metrics and hyperparameters used in these experiments are provided in Appendices C and D. In addition, although our primary target is RL tasks, we also test LRRL using SGD to minimize stationary objective functions to provide a simpler evaluation, as performance in RL is highly multi-factorial.
@@ 4.1 Comparing LRRL with Standard Learning
In our first experiment, we consider a set of $5$ learning rates. We set different configurations of LRRL using these values and compare them against the DQN agent reaching the best performance, in terms of maximum average return, among the possible single rates. More specifically, the learning rates considered are
@@
The whole set $\mathcal{K}(5)$ of rates is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) values, and the last one ($\mathcal{K}_{\text{sparse}}(3)$) taking the lowest/middle/highest values. All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations. The results in Figure 1 demonstrate that LRRL using the full set $\mathcal{K}(5)$ substantially outperforms the best-performing single learning rate in 2 out of 6 games (Seaquest and Breakout) while remaining competitive in the others. As shown in Figure 7 in Appendix B.1, the baseline DQN performance is highly sensitive to the learning rate choice, requiring extensive hyperparameter tuning to identify the optimal rate. In contrast, LRRL dynamically selects learning rates during training, avoiding the need for such tuning. Additionally, the $3$-arm bandit variants highlight that while LRRL can explore multiple learning rates within a single run, the specific set of rates available still impacts performance. Overall, these results illustrate the advantage of LRRL, which achieves similar or better performance than the best-tuned DQN baseline while eliminating the computational cost of manual learning rate tuning.
Figure 1: Comparison between LRRL and standard DQN learning: 5 variants of LRRL with different learning rate sets (provided in Appendix D) are tested again the DQN algorithm reaching best performance among the possible learning rates. The mean and one-half standard deviations over 5 runs are respectively represented by the curves and shaded areas.
Next, we illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves. Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$. In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses.
Figure 2: Systematic sampling of normalized learning rates and returns over the training steps using LRRL $\mathcal{K}(5)$ with Adam optimizer, through a single run. For each episode, we show the selected learning rate using different colors. Lower rates are increasingly selected over time.
@@ 4.2 Combining and Comparing Schedulers with LRRL
In the following experiments, LRRL is combined with learning rate schedulers. Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds). We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
@@
Figure 3 shows that LRRL combined with schedulers could once again substantially increase the final performance of 2 out of 6 tasks compared to using a fixed scheduler. In the remaining tasks, LRRL remains at least competitive, while for most of them, it shows a better jumpstart performance. The dashed black line represents the maximal average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by Andrychowicz et al. [4], who linearly decay the learning rate to 00.
In all the experiments conducted so far, we have set the update window $\kappa$ parameter to 1, which allows the observation of a full episode before updating the learning rate. In Appendix B.3, we address how this value could influence the performance of LRRL by allowing the observation of more interactions.
Figure 3: LRRL with arms that select learning rate schedulers: LRRL vs 3 individual schedulers ($d=1,2,3$). The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
4.3 Comparing RMSProp and Adam Optimizers
@@
Another widely used optimizer for deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism. Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum. Although the standard RMSProp does not feature momentum, we found that adding momentum to RMSProp can increase both the performance of DQN and LRRL, aligning with findings in the literature [30, 4]. In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN. Since the results were consistent in previous sections, we extend the experiments to other 3 Atari games.
@@
Figure 4 shows that LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN. Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance. Future work should investigate whether this slow convergence is due to some optimizer features such as the absence of bias correction in the first and second moment estimates.
@@
4.4 LRRL with a Distributional RL Algorithm
Lastly, we illustrate the versatility of LRRL by combining it with the distributional RL algorithm Implicit Quantile Networks (IQN) [11], which exhibits the best performance in many Atari games among the algorithms available in Dopamine. Adam is used as the optimizer. Figure 5 shows that LRRL outperforms standard learning with a single learning rate across all the tasks evaluated. More generally, LRRL can be seamlessly integrated with any RL algorithm that employs function approximators based on SGD, improving its performance and mitigating hyperparameter tuning.
Figure 5: Comparison between LRRL and standard IQN learning: IQN combined with LRRL using 3 different learning rates is tested against the 3 variants of standard IQN using each possible learning rate.
4.5 SGD and Stationary Non-convex Loss Landscapes
Although RL tasks are a challenging showcase to illustrate the effectiveness of our approach in the non-stationary case, there are many confounding factors to analyze their impact on the results, such as the optimizer adaptive routine, the environment stochasticity, and the reward sparseness. To disentangle these factors from our evaluation, in the following experiments we employ the SGD optimizer, which has a single scalar learning rate without adaptive scheme or momentum to find the global minima of $6$ non-convex loss landscapes. Since these functions represent a stationary case, we employ the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS) algorithm (see details in Appendix B.2), using as feedback the loss plus a small constant $\xi$, such that $f_{n}=\frac{1}{|\textrm{loss}(n)|+\xi}$, observed at each gradient step $n$ to update the arm’s probabilities.
Figure 6 illustrates the performance of LRRL against SGD with a single learning rate, starting from the same initial point and for the same amount of steps. The final steps are indicated by a star icon with a different color. The results show that LRRL, with its “all-in-one” strategy, efficiently navigates intricate landscapes. It is effective in combining higher learning rates to accelerate convergence and lower ones for more fine-grained optimization steps when needed, as evidenced in the loss landscapes depicted in Figures 6(a)6(b), and 6(c). In addition, Figures 6(d)6(e), and 6(f) illustrate that LRRL can be also effective in converging towards the best-performing learning rate within the arms, showcased by its steady convergence towards the global minima.
Figure 6: Comparison between LRRL and fixed learning rates in a stationary setting with SGD. LRRL with 3 learning rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 10 in Appendix B.4). LRRL is effective in either converging to the best performing rate or combining the set to achieve better results.
@@ 5 Conclusion
In this work, we introduced LRRL, a meta-learning approach for selecting optimizers’ learning rates on the fly. In any optimization task, the best-performing learning rate or rate scheduler is typically unknown, requiring multiple runs to test each candidate in a wasteful computational effort. LRRL mitigates this cost by employing a multi-armed bandit algorithm to explore a set of learning rates within a single run, significantly reducing the need for extensive hyperparameter tuning while achieving better or competitive results with minimal computational overhead. Using RL and non-convex optimization tasks, our empirical results show that combining LRRL with the Adam or SGD optimizers could enhance the performance, outperforming baselines and learning rate schedulers.
Future investigations could extend LRRL ideas to other hyperparameters, such as mini-batch size, which also plays a key role in convergence. Moreover, although LRRL selects rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [33, 19, 1].
Limitations. The primary limitation of LRRL lies in the need to define a set of arms for the MAB algorithm, as the choice of this set can directly influence the performance of the underlying learning algorithm. However, in computationally intensive settings such as RL tasks, LRRL offers a significant advantage: it enables the simultaneous evaluation of multiple learning rates within a single training run, eliminating the need to test each rate individually. By employing an “all-in-one” strategy, LRRL not only achieves competitive results while reducing the burden of hyperparameter optimization but also can outperform standard learning by combining multiple learning rates.
Broader Impact Statement
This work aims to contribute to the advancement of the field of Machine Learning. While we did not identify any direct negative social impacts associated with this research, we encourage future studies building on it to remain vigilant and critically evaluate potential ethical and societal implications.
@@ Acknowledgment
Appendix A Related Work
This section presents works closely related to ours. We first introduce approaches that use MAB for hyperparameter selection and later other methods that aim to adapt the learning rate over the training process.
Multi-armed bandit for (hyper)parameter selection.
Deep RL is known to be overly optimistic in the face of uncertainty [28], and many works have addressed this issue by proposing conservative policy updates [42, 14, 2]. However, when the agent can interact with the environment, this optimism might encourage exploration, potentially leading to the discovery of higher returns. Building on this idea, Moskovitz et al. [27] proposes Tactical Optimism and Pessimism (TOP), an approach that uses an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process. Although our work shares similarities with TOP, such as employing an adversarial bandit algorithm and using performance improvement as feedback, there are key differences. First, TOP is closely tied to actor-critic methods, making it unclear how it could be extended to other RL paradigms. Moreover, it also introduces an exploration strategy which depends on epistemic uncertainty (captured through ensembles) and aleatoric uncertainty (estimated using a distributional RL algorithm). Second, TOP trades-off between optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning. In contrast, LRRL offers an algorithm-agnostic solution by directly optimizing the learning rate, a core hyperparameter in all SGD-based methods, making it simpler and more broadly applicable. Also close to our work, Liu et al. [24] propose Adam with Bandit Sampling (ADAMBS), which employs the Exp3 algorithm to enhance sample efficiency by incorporating importance sampling within the Adam optimizer. ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 34).
In order to select the learning rate for stochastic gradient MCMC, Coullon et al. [10] employ an algorithm based on Successive Halving [17, 16], a MAB strategy that promotes promising arms and prunes suboptimal ones over time. In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [22, 36]. A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate on the fly from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space. Close to our work, Parker-Holder et al. [29] propose population-based training which employs a Gaussian Process (see details in 31) with a time-varying kernel to represent the upper confidence bound in a bandit setting. The idea is to have many instances of agents learning a task in parallel with different hyperparameters, replacing those that are underperforming within a single run.
Learning rate adapters/schedulers.
Optimizers such as RMSProp [41] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [18] also incorporates momentum to smooth gradient steps. However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 13). Increment-Delta-Bar-Delta (IDBD), introduced by Sutton [38], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [47]. Learning rate schedulers with time decay [35, 46] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning. A meta-gradient RL method is proposed in Xu et al. [45], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function. Our work differs from these methods by employing a MAB approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
Appendix B Supplementary Experiments
B.1 Baseline Evaluation with Varying Learning Rates
@@
To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer. The results presented in Figure 7 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to worse jumpstart performance and slow convergence.
@@
Figure 7: DQN performance using the Adam optimizer with varying learning rates across 6 Atari games. Shaded region represents one-half standard deviation. The results show that the best performing learning rate varies across games.
B.2 A Comparison between Multi-Armed Bandits Algorithms
@@
Adversarial MAB algorithms are designed for environments where the reward distribution changes over time. In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown probability distributions. To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5]. MOSS trades off exploration and exploitation by pulling the arm $k$ with highest upper confidence bound given by:
@@
where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to round $n$.
In Figure 8, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balances exploration and exploitation in MOSS. Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$. The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
Figure 8: A comparison between adversarial (Exp3) and stochastic (MOSS) MAB algorithms. Shaded region represents one-half standard deviation. The results show that Exp3 performs better overall.
B.3 Dealing with Noise and Lack of Feedback
LRRL relies entirely on performance improvements to update the arm probabilities. However, some RL environments can be highly stochastic, with sparse or delayed rewards, which increases the difficulty of learning tasks and, consequently, the ability of LRRL to make optimal choices. To mitigate this limitation, we introduce the variable $\kappa$ in Algorithm 1, which determines the number of interactions to observe before making a decision.
When setting $\kappa$, two factors should be considered:
  • The network’s update rate: Selecting new learning rates more frequently than updating the model leads to inaccurate feedback for the bandit mechanism since no model update might occur after pulling some arm.
  • The number of environment steps: In some tasks, positive or negative outcomes of a sequence of actions are only perceived at the end of an episode, and capturing sufficient information about the agent performance is essential to updating the arms’ probabilities.
In the experiments presented in this work, we set $\kappa=1$ episode, allowing the observation of a complete episode before updating the arms’ probabilities. To explore the impact of this parameter further, we conduct an ablation study using different $\kappa$ values to evaluate potential performance improvements when combined with the schedulers shown in Figure 3. As suggested by Figure 9, the choice of $\kappa$ is task-dependent: while some tasks benefit from more reactive learning rate updates, others show marginal improvements with increased observation counts. Ultimately, setting $\kappa=1$ seems relatively robust for Atari games, although higher values could slightly increase the performance in some tasks.
Figure 9: Results of the ablation study evaluating the impact of varying $\kappa$ with LRRL using schedulers. The dashed black line represents the max average return achieved by the best performing scheduler.
B.4 Supplementary Results using SGD
In Section 4.5, we tested the SGD optimizer in $6$ non-convex loss landscapes show in Figure 10. These functions were selected since they represent common benchmarks in optimization, each one offering different challenges due to their distinct shapes. By using SGD on these landscapes, we minimize the influence of confounding factors, such as the high stochasticity inherent in some RL tasks and the complexity of more robust optimizers like Adam.
Figure 11 presents the normalized loss and the arms selected at each optimization step. The results demonstrate LRRL’s dynamic adaptation mechanism, either converging to the most effective learning rate or combining them to achieve superior outcomes.
Figure 10: The 6 stationary non-convex loss landscapes used to assess LRRL with SGD optimizer.
Figure 11: Learning rates and logarithmic scale of losses over the gradient steps using LRRL with SGD optimizer.
Appendix C Evaluation Metrics
In this section, we describe the evaluation metrics that can be used to evaluate the agent’s performance as it interacts with an environment. Based on Dopamine, we use the evaluation step-size “iterations”, defined as a predetermined number of episodes. Figure 12 illustrates the evaluation metrics used in this work, as defined in Taylor and Stone [40]: